OpenStudy (anonymous):

A function of the form f(x) = mx + b, where m and b are real numbers, is called a _____ function. Example: f(x) = 6x - 5

OpenStudy (anonymous):

please help

Nnesha (nnesha):

f(x) same as y so y =mx+b <--- this is linear equation slope intercept so that function also same like this one y =mx+b is that linear or exponential function

OpenStudy (anonymous):

i dont know

OpenStudy (triciaal):

a function in the form f(x) = y = m x + b is a function with the highest exponent is one this give you a line therefore linear

Nnesha (nnesha):

y = mx+b <=== linear equation f(x) = mx+b <===this one also same they just wrote f(x) which represent function instead of y and f(x)=y they both are same equation if first one is linear then 2nd one also ????

OpenStudy (triciaal):

for example when the highest power is 2 then quadratic you have a parabola